# Token Interdependency Parsing (Tipping) - Fast and Accurate Log Parsing

### v1

@@ Abstract
In the last decade, an impressive increase in software adaptions has led to a surge in log data production, making manual log analysis impractical and establishing the necessity for automated methods. Conversely, most automated analysis tools include a component designed to separate log templates from their parameters, commonly referred to as a ”log parser”.
@@
Index Terms:
Log Parser, Strongly Connected Components, Automated Log Analysis, Large Scale Log Parsing, Parallel Processing
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ 1 Introduction
In recent years, breakthroughs in software development have led to an exponential increase in log generation, as complex systems and applications continuously produce vast amounts of log data. These logs provide insights into system performance, user behaviors, and potential security threats. However, the sheer volume and speed at which this data is produced have made it increasingly challenging for human analysts to manually process and interpret these logs effectively. As software systems grow more sophisticated and interconnected, the deluge of data, including transaction details, system statuses, and operational anomalies, continues to expand.
@@
The log parser is one of the most crucial components in any log analysis pipeline. This tool is the foundational mechanism that interprets raw log messages into a structured format, making it crusial for further analysis. Log parsing, sometimes referred to as log clustering, is a method aimed at simplifying the complexity of free-form log messages. This process involves differentiating between a log message’s constant and variable parts; the constant part (template) represents the event description, while the variable part (parameters) includes dynamic data specific to each event. As an illustration, in the log message ”Socket connected successfully to 10.10.1.10”, the phrase ”Socket connected successfully to” is the template, which indicates an event. At the same time, ”10.10.1.10” represents the replaceable parameter that may vary with each specific event occurrence.
@@
Furthermore, log parsing plays a vital role in enhancing software management and security across various fields[1]. It is essential for anomaly detection [2, 3, 4, 5, 6], aiding in the identification and prediction of unusual behaviors through advanced algorithms. For debugging [7, 8], it enables developers to trace errors back to specific lines of code or external factors. Log parsing is used to identify bottlenecks, improving application efficiency and speed in performance analysis [9, 10]. Security analysis [11, 12] leverages log parsing to detect potential threats by analyzing patterns of suspicious activity. Real-time software monitoring [13] utilizes parsed logs to gain insights into user behavior and system performance, facilitating issue detection. Root cause analysis [14, 15] uses log parsing to model common event patterns, helping to identify the origins of software issues. Log parsing supports audit trails [16] by tracking data access and modifications, ensuring compliance and security. Each of these applications underscores the versatility and indispensability of log parsing in modern software engineering. The pivotal importance of log parsing has prompted the creation of numerous tools aimed at optimizing this process [17, 18, 19, 20, 21, 22, 23, 24]
@@
In log parsing, performance and accuracy stand as the twin pillars essential to its effectiveness. Performance refers to the speed at which log data can be processed and analyzed, a critical factor when dealing with large volumes of logs. Accuracy, however, involves the precision with which the log parser can identify and separate the template and parameter parts from each other, which is crucial for meaningful interpretation and making the most of down-the-stream components. However, these two often find themselves at odds. Enhancing accuracy typically involves more complex processing techniques that are computationally intensive and slower. This inverse relationship poses a significant challenge: striving for higher accuracy can compromise the speed.
@@
Parallel processing refers to distributing tasks across multiple processors or cores to execute operations simultaneously, enhancing speed as a result. In the context of log parsing, this approach enables the handling of extensive datasets at a faster rate while allowing for the incorporation of complex parsing algorithms that ensure high accuracy. Log parsing performance can be improved by leveraging parallel processing while maintaining accuracy as complicated methods are broken down into chunks that can be processed concurrently.
However, implementing parallel processing comes with its own set of challenges. Managing concurrent operations across different processors adds complexity to system design, necessitating mechanisms for coordinating tasks and handling potential conflicts, such as mutual exclusion. Additionally, ensuring data consistency and synchronization across multiple processing streams can be difficult, posing a risk of errors and inconsistencies in the parsed results. Thus, while parallel processing offers substantial benefits in terms of performance and capability for sophisticated methods, it requires meticulous calculations and robust architecture design to overcome these challenges.
@@
Finally, we demonstrate that the Tipping algorithm outperforms state-of-the-art methods across four benchmarks, highlighting its accuracy and performance. Tipping achieves higher accuracy and outputs fast parsing times in the LogPai and LogPM tests for accuracy and performance and processes a massive volume of logs faster than other methods in the LogLead test for computational efficiency. Additionally, Tipping excels in LogLead’s anomaly detection downstream task as well. These results underscore Tipping’s robustness and adaptability, making it a performance, accurate, and versatile tool for various log analysis applications.
@@ 2 Related works
In a brilliant study, Drain [18] is designed to process logs in a streaming and timely manner. Unlike traditional log parsers focusing on offline batch processing and struggling with the rapidly increasing volume of logs, Drain introduced a fixed depth parse tree to accelerate the parsing process. This parse tree, crucial for guiding the log group search process, encodes specially designed parsing rules within its nodes. Its fixed depth tree structure and specially designed parsing rules significantly enhance the speed and accuracy of log parsing, making it an excellent tool for online log parsing.
@@
Another study introduces the Brain[17], an innovative log parsing approach that leverages a bidirectional parallel tree to transform semi-structured logs into a structured format, aimed at overcoming the limitations of existing parsers in software systems. By identifying the longest common pattern across logs as a likely component of the log template, Brain efficiently groups logs and hierarchically integrates constant words into these patterns to form complete templates. This method significantly enhances parsing accuracy and efficiency across various software logs. Furthermore, Brain’s process encompasses preprocessing for word splitting and variable filtering, initial group creation based on common patterns, and using a bidirectional tree to systematically add constant words, culminating in the generation of structured log templates. Brain offers improved stability and processing speed for large volumes of log data.
@@
NuLog, introduced in [25], is a self-supervised log parsing technique leveraging transformer architecture to address the challenge of parsing massive volumes of semi-structured logs generated by large-scale software systems. Unlike traditional methods that rely on heuristics or manual rule extraction, NuLog uses masked language modeling (MLM) to predict the presence of words in log messages based on their context, distinguishing between constant and variable parts without requiring domain knowledge. The methodology encompasses tokenization, masking, and a multi-head self-attention mechanism with positional encoding and a feed-forward network, outputting log templates alongside numerical vector summarizations. This enables high parsing accuracy and supports downstream anomaly detection tasks in both supervised and unsupervised scenarios. NuLog’s approach significantly advances automated log analysis, offering broad applicability and minimal human intervention.
@@
Another ingenious paper presents Spell[19], a novel streaming log parser that leverages the Longest Common Subsequence (LCS) method for parsing system event logs in real time into structured data. Moreover, Spell dynamically extracts log patterns and updates its log structure by comparing incoming log entries against stored LCS sequences within an LCSMap data structure. Furthermore, It utilizes a pre-filtering step using a prefix tree and a simple loop approach to efficiently identify existing message types, significantly reducing the computational load by applying LCS computation selectively. Spell is considered an effective solution for online log monitoring and analysis, offering immediate insights into system behavior and enhancing log management capabilities.
@@
MoLFI, introduced in [20], proposes an innovative approach by framing it as a multi-objective optimization problem, efficiently solved using the NSGA-II [26] algorithm to balance the frequency and specificity of templates. This method requires no domain knowledge to address the challenges of rapidly evolving log formats. MoLFI produces a Pareto optimal set of templates that adeptly navigates the trade-offs between template commonality and uniqueness through multiple steps of preprocessing, solution encoding, and iterative refinement using genetic operations. Overall, MoLFI introduced a completely different approach compared to other well-known log parsers at the time.
@@
Logram[21] is another log parsing method that employs n-gram dictionaries to address the log parsing problem efficiently. By leveraging the intuition that frequently occurring n-grams likely represent static text and rare n-grams indicate dynamic variables, Logram generates n-gram dictionaries from log messages and uses these dictionaries to parse and distinguish between static and dynamic components of logs. This process involves preprocessing logs to extract tokens, creating dictionaries for efficient queries, and employing an automated method to determine the occurrence threshold for dynamic variable identification. In their own evaluations, Logram showed improved parsing accuracy and efficiency.
@@
ULP (Unified Log Parser)[22] is a log parsing tool that enhances accuracy and efficiency through a novel combination of string matching and local frequency analysis. ULP effectively differentiates between static and dynamic components of log messages by grouping log events based on string similarity and then applying frequency analysis within these groups. This method outperformed its rival competition at the time. The parser involves preprocessing to identify trivial dynamic tokens, grouping similar events for targeted analysis, and generating templates through local frequency insights, showcasing a significant improvement in parsing large log files and offering a solution for complex log analysis tasks.
@@
The log parser LenMa (Length Matters) [27] introduces a unique clustering approach, leveraging the length of words within each message to distinguish between fixed and variable components, thereby facilitating efficient online template generation. LenMa simplifies the process by employing word length vectors to compute similarity scores between new log entries and existing clusters using cosine similarity. This innovative approach enables real-time, dynamic clustering of log messages, efficiently adapting to the continuous influx of logs. LenMa proves to be an effective and straightforward method for managing vast volumes of log data in a scalable solution.
@@
IpLom (Iterative Partitioning Log Mining) [28, 29] is a multi-step log clustering approach that hierarchically partitions log messages into different groups. The first step of partitioning divides log messages into groups based on the number of tokens. The second step further processes each of these groups by splitting them based on the token position with the least variability, i.e., they are the most likely to be templates. The third step searches for a bijective relationship with token positions that co-vary. Although presented already 15 years ago, IpLom has shown promising accuracy and speed in independent tests [1]. The main drawback of IpLom is that varying numbers of tokens in log parameters can cause incorrect groups, and the algorithm does not include the ability to correct these groups.
@@
AEL (Abstracting Execution Logs) [30] is based on four hierarchical steps: anonymize, tokenize, categorize, and reconcile. The anonymized step uses external heuristics to convert log message parameters to generic tokens. The tokenize step uses the number of tokens and parameters from the previous step to group log messages. The categorization step goes through existing groups and performs token-by-token comparison and clustering. Finally, the reconcile step tries to account for the lack of heuristics of the anonymized step, and merges log clusters if too many execution events have been created. AEL has also shown good accuracy and speed in independent tests [1]. The main drawback of the algorithm is the need to have A-priori heuristics on how to recognize log message parameters.
@@ 3 Proposed Method
This section outlines the structure of our methodology, breaking it down into its components. Our algorithm is structured into two phases, unlike other methods that perform parsing in a single phase. During the initial phase (Sec. 3.1 to 3.3), it gathers information from log messages. In the subsequent phase (Sec. 3.4 to 3.10), the collected information is employed for log parsing. The two-phase approach enhances the algorithm’s accuracy and allows for the utilization of parallel processing.
Figure 1 shows an overview of our approach, while the following subsections explain each phase in detail.
@@ 3.1 Tokenizer
The first component of our parser is the tokenizer, which is common in almost every NLP task. The tokenizer breaks down log messages into smaller substrings that make the data more suitable for algorithms to process. Unlike most previous methods that split text at white space locations, our approach to tokenization allows for domain knowledge to be passed from users.
@@
The white special patterns refer to an optional list of regular expressions that the algorithm should recognize as templates during the parsing process. These patterns are earmarked to be excluded from being recognized as parameters, ensuring they remain intact within the entire parsing process and will become a part of the output templates.
Conversely, black special patterns are designated to be recognized as parameters unequivocally, ensuring their segregation from the template content. White and black special patterns allow for more nuances in the parsing process and results, resolving problems such as maintaining template-shared tokens[1] and execution environment bias parameters[1].
Lastly, the symbols input consists of a curated set of punctuation characters. These symbols, alongside whitespace characters, serve as delimiters in the tokenization process to define the boundaries between non-special tokens. Our algorithm achieves a sophisticated and tailored approach toward log message tokenization by leveraging these three inputs.
The tokenizer initiates its process by scanning the log message to identify all occurrences of white special patterns, which are then marked as white tokens. Subsequently, the algorithm examines the remaining portions of the log message to locate black special patterns and flag them as black tokens. Following the identification of white and black tokens, the algorithm continues with an iteration over the residual segments of the log message while employing symbols and white space characters as delimiters to fragment the message. Uniquely, while splitting the message, the algorithm retains these delimiter characters as separate tokens.
After identifying white and black tokens, the remainder of the log message is fragmented by symbols and whitespaces. The tokenizer concludes its process by compiling the sequence of tokens and returning it as the tokenized message.
@@ 3.2 Token Filter
The token filter receives tokenized messages and filters out unwanted tokens based on user-defined parameters. This allows for better leverage of domain knowledge while being capable of producing accurate results using the default values. By default, the token filter only keeps the white special tokens and purely alphabetic tokens.
@@ 3.3 Occurences and Co-occurrences
For further down-the-line components, keeping track of both token occurrences and their co-occurrences is essential. Token occurrence refers to the number of times that a particular token appears within a given dataset. On the other hand, co-occurrence involves tracking instances where two tokens appear together within the same message. Both occurrence and co-occurrence could be stored in a single sparse matrix data structure. However, storing occurrences and co-occurrences in a hash map data structure is more suitable, as row-or-column-wise access is not required. Within the hash map data structure, each key is a single-or-two-members set of token(s), indicating the occurrences and co-occurrences, while the value is the count for that set.

### v2

# Token Interdependency Parsing (Tipping) - A Fast and Accurate Statistical Log Parser

@@ Abstract
In the last decade, an impressive increase in software adaptations has led to a surge in log data production, making manual log analysis impractical and establishing the necessity for automated methods. Conversely, most automated analysis tools include a component designed to separate log templates from their parameters, commonly referred to as a ”log parser”.
@@
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
@@ 1 Introduction
In recent years, breakthroughs in software development have led to an exponential increase in log generation, as complex systems and applications continuously produce vast amounts of log data. These logs provide insights into system performance, user behaviours, and potential security threats. However, the sheer volume and speed at which this data is produced have made it increasingly challenging for human analysts to manually process and interpret these logs effectively. As software systems grow more sophisticated and interconnected, the deluge of data, including transaction details, system statuses, and operational anomalies, continues to expand.
@@
The log parser is one of the most crucial components in any log analysis pipeline. This tool is the foundational mechanism that interprets raw log messages into a structured format, making it crucial for further analysis. Log parsing, sometimes referred to as log clustering, is a method aimed at simplifying the complexity of free-form log messages. This process involves differentiating between a log message’s constant and variable parts; the constant part (template) represents the event description, while the variable part (parameters) includes dynamic data specific to each event. As an illustration, in the log message ”Socket connected successfully to 10.10.1.10”, the phrase ”Socket connected successfully to” is the template, which indicates an event. At the same time, ”10.10.1.10” represents the replaceable parameter that may vary with each specific event occurrence.
@@
Furthermore, log parsing plays a vital role in enhancing software management and security across various fields[logpm]. It is essential for anomaly detection [sialog, loganomaly, deeplog, cnnlog, monika], aiding in the identification and prediction of unusual behaviors through advanced algorithms. For debugging [logdebugging1, logdebugging2], it enables developers to trace errors back to specific lines of code or external factors. Log parsing is used to identify bottlenecks, improving application efficiency and speed in performance analysis [netlogger, loggingperformance]. Security analysis [securitylog1, securitylog2] leverages log parsing to detect potential threats by analyzing patterns of suspicious activity. Real-time software monitoring [logsbusiness] utilizes parsed logs to gain insights into user behaviour and system performance, facilitating issue detection. Root cause analysis [rcalog1, rcalog2] uses log parsing to model common event patterns, helping to identify the origins of software issues. Log parsing supports audit trails [audit] by tracking data access and modifications, ensuring compliance and security. Each of these applications underscores the versatility and indispensability of log parsing in modern software engineering. The pivotal importance of log parsing has prompted the creation of numerous tools aimed at optimizing this process [brain, drain, spell, molfi, logram, ulp, loghub, loghubtools]
@@
In log parsing, performance and accuracy stand as the twin pillars essential to its effectiveness. Performance refers to the speed at which log data can be processed and analyzed, a critical factor when dealing with large volumes of logs. Accuracy, however, involves the precision with which the log parser can identify and separate the template and parameter parts from each other, which is crucial for meaningful interpretation and making the most of downstream components. However, these two often find themselves at odds. Enhancing accuracy typically involves more complex processing techniques that are computationally intensive and slower. This inverse relationship poses a significant challenge: striving for higher accuracy can compromise the speed.
@@
Parallel processing refers to distributing tasks across multiple processors or cores to be executed simultaneously, enhancing speed as a result. Accordingly, log parsing can be improved by leveraging parallel processing, as parallel processing allows for handling extensive datasets at a faster rate while allowing for the incorporation of complex parsing algorithms that ensure high accuracy. However, implementing parallel processing comes with its own set of challenges, such as complexity, mutual exclusion, and ensuring data consistency and synchronization.
@@
Finally, we demonstrate that the Tipping algorithm outperforms state-of-the-art methods across four benchmarks, highlighting its accuracy and performance. Tipping yields the fastest parsing times in the LogHub2k, LogHub2.0, and LogPM benchmarks while maintaining accuracy, and processes a massive volume of logs faster than other methods in the LogLead for computational efficiency. Additionally, Tipping excels in LogLead’s anomaly detection downstream task as well. These results underscore Tipping’s robustness and adaptability, making it a performant, accurate, and versatile tool for various log analysis applications.
@@ 2 Related works
In a brilliant study, Drain [drain] is designed to process logs in a streaming and timely manner. Unlike traditional log parsers focusing on offline batch processing and struggling with the rapidly increasing volume of logs, Drain introduced a fixed-depth parse tree to accelerate the parsing process. This parse tree, crucial for guiding the log group search process, encodes specially designed parsing rules within its nodes. Its fixed-depth tree structure and specially designed parsing rules significantly enhance the speed and accuracy of log parsing, making it an excellent tool for online log parsing.
@@
Another study introduces the Brain[brain], an innovative log parsing approach that leverages a bidirectional parallel tree to transform semi-structured logs into a structured format, aimed at overcoming the limitations of existing parsers in software systems. By identifying the longest common pattern across logs as a likely component of the log template, Brain efficiently groups logs and hierarchically integrates constant words into these patterns to form complete templates. This method significantly enhances parsing accuracy and efficiency across various software logs. Furthermore, Brain’s process encompasses preprocessing for word splitting and variable filtering, initial group creation based on common patterns, and using a bidirectional tree to systematically add constant words, culminating in the generation of structured log templates. Brain offers improved stability and processing speed for large volumes of log data.
@@
NuLog, introduced in [nulog], is a self-supervised log parsing technique leveraging transformer architecture to address the challenge of parsing massive volumes of semi-structured logs generated by large-scale software systems. Unlike traditional methods that rely on heuristics or manual rule extraction, NuLog uses masked language modeling (MLM) to predict the presence of words in log messages based on their context, distinguishing between constant and variable parts without requiring domain knowledge. The methodology encompasses tokenization, masking, and a multi-head self-attention mechanism with positional encoding and a feed-forward network, outputting log templates alongside numerical vector summarizations. This enables high parsing accuracy and supports downstream anomaly detection tasks in both supervised and unsupervised scenarios. NuLog’s approach significantly advances automated log analysis, offering broad applicability and minimal human intervention.
@@
Another ingenious paper presents Spell[spell], a novel streaming log parser that leverages the Longest Common Subsequence (LCS) method for parsing system event logs in real-time into structured data. Moreover, Spell dynamically extracts log patterns and updates its log structure by comparing incoming log entries against stored LCS sequences within an LCSMap data structure. Furthermore, It utilizes a pre-filtering step using a prefix tree and a simple loop approach to efficiently identify existing message types, significantly reducing the computational load by applying LCS computation selectively. Spell is considered an effective solution for online log monitoring and analysis, offering immediate insights into system behavior and enhancing log management capabilities.
@@
MoLFI, introduced in [molfi], proposes an innovative approach by framing it as a multi-objective optimization problem, efficiently solved using the NSGA-II [nsgaii] algorithm to balance the frequency and specificity of templates. This method requires no domain knowledge to address the challenges of rapidly evolving log formats. MoLFI produces a Pareto optimal set of templates that adeptly navigates the trade-offs between template commonality and uniqueness through multiple steps of preprocessing, solution encoding, and iterative refinement using genetic operations. Overall, MoLFI introduced a completely different approach compared to other well-known log parsers at the time.
@@
Logram[logram] is another log parsing method that employs n-gram dictionaries to address the log parsing problem efficiently. By leveraging the intuition that frequently occurring n-grams likely represent static text and rare n-grams indicate dynamic variables, Logram generates n-gram dictionaries from log messages and uses these dictionaries to parse and distinguish between static and dynamic components of logs. This process involves preprocessing logs to extract tokens, creating dictionaries for efficient queries, and employing an automated method to determine the occurrence threshold for dynamic variable identification. In their own evaluations, Logram showed improved parsing accuracy and efficiency.
@@
ULP (Unified Log Parser)[ulp] is a log parsing tool that enhances accuracy and efficiency through a novel combination of string matching and local frequency analysis. ULP effectively differentiates between static and dynamic components of log messages by grouping log events based on string similarity and then applying frequency analysis within these groups. This method outperformed its rival competition at the time. The parser involves preprocessing to identify trivial dynamic tokens, grouping similar events for targeted analysis, and generating templates through local frequency insights, showcasing a significant improvement in parsing large log files and offering a solution for complex log analysis tasks.
@@
The log parser LenMa (Length Matters) [lenma] introduces a unique clustering approach, leveraging the length of words within each message to distinguish between fixed and variable components, thereby facilitating efficient online template generation. LenMa simplifies the process by employing word length vectors to compute similarity scores between new log entries and existing clusters using cosine similarity. This innovative approach enables real-time, dynamic clustering of log messages, efficiently adapting to the continuous influx of logs. LenMa proves to be an effective and straightforward method for managing vast volumes of log data in a scalable solution.
@@
IpLom (Iterative Partitioning Log Mining) [makanju2009clustering, makanju2011lightweight] is a multi-step log clustering approach that hierarchically partitions log messages into different groups. The first step of partitioning divides log messages into groups based on the number of tokens. The second step further processes each of these groups by splitting them based on the token position with the least variability, i.e., they are the most likely to be templates. The third step searches for a bijective relationship with token positions that co-vary. Although presented already 15 years ago, IpLom has shown promising accuracy and speed in independent tests [logpm]. The main drawback of IpLom is that varying numbers of tokens in log parameters can cause incorrect groups, and the algorithm does not include the ability to correct these groups.
@@
AEL (Abstracting Execution Logs) [jiang2008automated] is based on four hierarchical steps: anonymize, tokenize, categorize, and reconcile. The anonymized step uses external heuristics to convert log message parameters to generic tokens. The tokenize step uses the number of tokens and parameters from the previous step to group log messages. The categorization step goes through existing groups and performs a token-by-token comparison and clustering. Finally, the reconcile step tries to account for the lack of heuristics of the anonymized step and merges log clusters if too many execution events have been created. AEL has also shown good accuracy and speed in independent tests [logpm]. The main drawback of the algorithm is the need to have A-priori heuristics on how to recognize log message parameters.
@@ 3 Proposed Method
Tipping is divided into two phases. During the initial phase (Sec. 3.1 to 3.3), it gathers information from log messages. In the subsequent phase (Sec. 3.4 to 3.10), the collected information is employed to cluster messages, find templates, and generate parameter masks. This two-phase approach enhances the algorithm’s accuracy and allows for parallel processing.
Figure 1 visualize an overview of our approach, while the following subsections explain each phase in detail. The tokenizer, which divides log messages into smaller parts, is discussed at first. Then, interdependency graph construction and anchor token identification are described. The final steps involve clustering messages based on these anchor tokens and generating event templates and parameter masks.
@@ 3.1 Tokenizer
The first component of our parser is the tokenizer, which is common in almost every NLP task. The tokenizer breaks down log messages into smaller substrings that make the data more suitable for algorithms to process. Additionally, our approach allows for domain knowledge to be passed from users to the tokenizer using regular expressions.
@@
The white special patterns are a list of regular expressions that should always be recognized as templates. Conversely, black special patterns should be recognized as parameters.
Special patterns allow for more nuanced control in the parsing process, resolving problems such as maintaining template-shared tokens[logpm] and execution environment bias parameters[logpm]. Lastly, the symbols consist of a set of punctuation characters that, alongside whitespace characters, serve as delimiters in the tokenization process.
The tokenizer initiates its process by scanning the log message to identify and mark all occurrences of white special patterns. Subsequently, the algorithm examines the remaining parts to locate and mark black special patterns. Then, the algorithm iterates over the rest of the segments and uses symbols as delimiters to split the message.
@@ 3.2 Token Filter
The token filter receives tokenized messages and filters out tokens based on user-defined rules. This not only allows for better leverage of domain knowledge but also results in fewer tokens and less processing, as the results for the next components. By default, the token filter only keeps the white special tokens and purely alphabetic tokens.
@@ 3.3 Occurences and Co-occurrences
For further next components, keeping track of both token occurrences and their co-occurrences is essential.
Token occurrence refers to the number of times that a particular token appears within a given dataset. On the other hand, co-occurrence involves tracking instances where two tokens appear together within the same message. Both occurrence and co-occurrence could be stored in a single sparse matrix data structure. However, a hash map is faster and more memory efficient, as row- or column-wise access is not required. Within the hash map data structure, keys are single or two-member token sets, indicating occurrences or co-occurrences, while the value is the count for the keys.